# Bell Nonlocality from Wigner Negativity in Qudit Systems

### v2

@@ Abstract
The details of the evaluation are given in Eq. (22).
Measuring a complete set of $\{A_{u},\,u\in\mathbb{Z}_{d}^{2}\}$ restricts the lhv model, which needs to comply with the operators’ algebra. The operators $A_{u}$ are projectors and have an equivalent description in terms of the Pauli operators $T_{u}$ and vice versa by means of the Fourier transform. Given a set of measurements described by Pauli operators whose eigenbases are mutually unbiased, their outcomes are multiplicative characters of the form $\omega^{[a,u]}$ for $a\in\mathbb{Z}_{d}$ due to the structure of the Pauli group, specifically its intersection into subgroups $\mathcal{G}_{r}$ for $r=0,\dots,d$. As a result, a deterministic classical value assignment (lhv) can only assign deterministic values 00 and $d$ to the operator’s outcome, $d\left\langle A_{u}\right\rangle_{\mathrm{lhv}^{\ast}}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]-[a,v]}=d^{2}\delta_{u,a}$ for $a\in\mathbb{Z}_{d}^{2}$ and, hence, $\left\langle\mathcal{B}_{\nu}\right\rangle_{\mathrm{lhv}^{\ast}}=d^{2}\,W^{\nu}_{a_{1},a_{2}}\,.$ As a result, any lhv model, a convex mixture of deterministic classical value assignment (lhv), can maximally achieve
$\left\langle\mathcal{B}_{\nu}\right\rangle_{\mathrm{lhv}}\leq d^{2}\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}:=B_{\mathrm{lhv}}^{\max}\,.$ (13)
Since $|W^{\nu}_{v_{1},v_{2}}|\leq 1/d^{2}$, we obtain $B_{\mathrm{lhv}}^{\max}\leq 1=\langle\Phi|\mathcal{B}_{\nu}|\Phi\rangle$. On the other hand, from Eq. (25), the expectation value with a quantum state $\sigma$ can be bound by $B_{\mathrm{lhv}}^{\max}$ and the volume of Wigner negativity using Hölder’s inequality with the 1-norm and the maximum norm,
$\mathrm{Tr}\left(\mathcal{B}_{\nu}\sigma\right)\leq B_{\mathrm{lhv}}^{\max}(1+2N[(U^{\dagger}_{\nu}\otimes\mathds{1})\sigma(U_{\nu}^{\phantom{\dagger}}\otimes\mathds{1})])\,,$ (14)
and in particular,
$1\leq B_{\mathrm{lhv}}^{\max}\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}\left|W_{v_{1},v_{2}}^{\nu}\right|=B_{\mathrm{lhv}}^{\max}(1+2N[\Phi_{\nu}])\,.$ (15)
Thus, a Bell violation $\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=1>B_{\mathrm{lhv}}^{\max}$ is possible for a nontrivial negativity volume, $N[\Phi_{\nu}]>0$. The previous analysis of character polynomials defining the Wigner function $W^{\nu}_{v_{1},v_{2}}$, in particular Eq. (7), gives $B_{\mathrm{lhv}}^{\max}\leq 2/\sqrt{d}$ for $d>3$. For $d=3$, an exhaustive search using Eq. (17) finds $B_{\mathrm{lhv}}^{\max}<0.845$. As a result, we indeed obtain a Bell violation $1>d^{2}\max W^{\nu}_{v_{1},v_{2}}=B_{\mathrm{lhv}}^{\max}$. Table 1 lists such values $d^{2}\max W^{\nu}_{v_{1},v_{2}}$ for $d\leq 23$.
For a more compact Bell operator, we now focus solely on the stabilizer elements, $T_{(x,z)}\otimes T_{(x,-z)}|\Phi\rangle=|\Phi\rangle$, which reduces the number of measurements from $(d+1)^{2}$ to $(d+1)$, and leads to
@@
The expectation values have the same form as for the full Bell operator, but the summation and the maximum take only coefficients $x,z,t\in\mathbb{Z}_{d}$. The Bell operator is a measure of the singlet fraction and the volume of Wigner negativity. The Bell inequality is structurally the same and Bell state achieves the same Bell violation.
@@
where $X_{(1/3)}=UXU^{\dagger}$, and $U=\mathrm{diag}(1,\omega^{2/3},\omega^{1/3})$. Note that these operators $O=X,X_{(1/3)}$ are not Hermitian but unitary yet represent valid measurements in terms of their Fourier transform with projectors $\Pi_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}O^{k}$ of outcome $m\in\mathbb{Z}_{d}$. The same holds for a deterministic value assignment. For operators that have the same algebra as the Pauli operators $T_{u}$, it resorts to multiplicative characters of the form $\omega^{[a,u]}$ for $a\in\mathbb{Z}_{d}$, since the results of two commuting operators are equal when measured sequentially. As a result, any lhv model attains $B_{3,\mathrm{lhv}}^{\max}=\max(\omega^{a_{0}+b_{0}}+\omega^{2a_{0}+2b_{1}}+\omega^{2a_{1}+2b_{0}}+\omega^{a_{1}+b_{1}+1}+\,h.c.)=6+\omega+\omega^{2}=5$, where the maximum is taken over $a_{0},b_{0},a_{1},b_{1}\in\mathbb{Z}_{3}$, and is attained for $b_{1}=2$ and $a_{0},b_{0},a_{1}=0$. On the other hand, the quantum expectation value, $\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)=1+3(2\omega^{1/3}+\omega^{-2/3})+\,h.c.\,\approx 5.412\,,$ exceeds that of the classical model, $B_{3,\mathrm{lhv}}^{\max}<\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)$.
@@
$\mathcal{B}_{(1/2)}$ $=\frac{1}{d}\sum_{k\in\mathbb{Z}_{d}}X^{k}_{(1/2)}\otimes\left(X^{k}_{(1/2)}+\omega^{-k}X^{k}_{(-1/2)}\right)$
$+X^{k}_{(-1/2)}\otimes\left(X^{k}_{(-1/2)}+\omega^{k}X^{k}_{(1/2)}\right)\,.$ (19)
@@
The operator $\mathcal{B}_{S}$ is a measure for the overlap with the given stabilizer state under the condition that $((0,\mathbb{Z}_{d})_{1}\otimes(0,0)^{\otimes n-1})\subset\Sigma$, which is exactly the case if $|S\rangle$ is entangled over the cut of the first qudit. Then, $\langle S|\mathcal{B}_{S}|S\rangle=1$, while $\langle\mathcal{B}_{S}\rangle_{\mathrm{lhv}}\leq d^{n}\max_{{\bm{u}}\in\Sigma}W_{{\bm{u}}}^{\nu}<1$, with the same arguments as for the bipartite case. A single Bell operator of the form (20) cannot detect genuine multipartite entanglement GUHNE20091 . A family of Bell operators, where the unitary cube operator $U_{\nu}$ acts on a different qudit, can detect if all qudits are entangled with any other qudit, but fails to expose genuine multipartite entanglement.
@@
We have constructed a family of qudit Bell inequalities for any odd prime dimension. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $T$-gate characterized in Howard2012 . The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. The maximally entangled Bell state violates the corresponding Bell inequality and attains the quantum bound due to the separation between the 1-norm and the maximum norm of its rotated Wigner function. The Bell operator measures the singlet fraction and bounds the extent of Wigner negativity. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. Furthermore, the measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models. Moreover, nonlocal correlations can emerge with qudit Pauli operators when rotated by unitary operators with a non-character spectrum in bipartite systems. These operators have an obvious advantage, which is resorting to phases $\omega^{q}$ that are beyond the description of any local value assignment that can only resort to characters $\omega^{k}$. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving GKP (Gottesman–Kitaev–Preskill) states PhysRevA.64.012310 .
@@ Appendix A Wigner Negativity
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})}$
$\hskip 142.26378pt\cdot\langle jj|U^{\phantom{\dagger}}_{{f}}\,X^{(v_{1})_{x}}U^{\dagger}_{{f}}\otimes X^{(v_{2})_{x}}|kk\rangle$
@@
with the Pauli operators $\sigma_{(0,0)}=\mathds{1}$, $\sigma_{(1,0)}=X$, $\sigma_{(1,1)}=Y$, $\sigma_{(0,1)}=Z$. Note that for qubits the Bell operators in Eq. (10) and Eq. (16) do not lead to violation since any LHV model can assign independent variables to the qubit Pauli operators.
@@
$\mathrm{Tr}\left(\sigma\mathcal{B}\right)$ $=d^{2}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{u_{1},u_{2}}W_{u_{1},u_{2}}\left((U_{\nu})^{\dagger}\sigma\,U^{\phantom{\dagger}}_{\nu}\right)$
$\leq d^{2}\max_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{u_{1},u_{2}}|\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W_{u_{1},u_{2}}\left((U_{\nu})^{\dagger}\sigma\,U^{\phantom{\dagger}}_{\nu}\right)|$
$=B_{\mathrm{lhv}}^{\max}(1+2N[(U_{\nu})^{\dagger}\sigma\,U^{\phantom{\dagger}}_{\nu}])\,.$ (25)
Appendix B Bell Inequality for Stabilizer States

### v3

@@ Abstract
From Eq. (32), the expectation value with a quantum state $\sigma$ can be bound by $B_{\mathrm{nc}}^{\max}$ and the volume of Wigner negativity using Hölder’s inequality with the 1-norm and the maximum norm,
$\mathrm{Tr}\left(\mathcal{B}_{\nu}\sigma\right)\leq B_{\mathrm{nc}}^{\max}(1+2N[\sigma_{\nu}])\,,$ (18)
with $\sigma_{\nu}:=(U_{\nu}\otimes\mathds{1})^{\dagger}\sigma(U_{\nu}\otimes\mathds{1})$. In particular, it is $1\leq B_{\mathrm{nc}}^{\max}(1+2N[\Phi_{\nu}])\,.$ Thus, a non-contextual violation $\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=1>B_{\mathrm{lhv}}^{\max}$ is possible for a nontrivial negativity volume, $N[\Phi_{\nu}]>0$. The previous analysis of character polynomials defining the Wigner function $W^{\nu}_{v_{1},v_{2}}$, in particular Eq. (8), gives
$B_{\mathrm{nc}}^{\max}\leq 2/\sqrt{d}$ (19)
for $d>3$. For $d=3$, an exhaustive search using Eq. (21) finds $B_{\mathrm{nc}}^{\max}<0.844$. Table 1 lists such values $d^{2}\max W^{\nu}_{v_{1},v_{2}}[\Phi_{\nu}]$ for $d\leq 23$. For a classical model, the correlations from $\mathcal{B}_{\nu}$ are equivalent to those from $\mathcal{B}[\Phi_{\nu}]$. The maximum value attainable by an lhv model is the same and an exhaustive search finds tight values for $B^{\max}_{\mathrm{lhv}}$ for small odd prime $d$ in Tab. 1, which seem to decrease with growing $d$.
For a more compact Bell operator, we now focus solely on the stabilizer elements, $T_{(x,z)}\otimes T_{(x,-z)}|\Phi\rangle=|\Phi\rangle$, which reduces the number of measurements from $(d+1)^{2}$ to $d^{2}+1$, and leads to
@@
For convenience, we use the operator $A_{u}$, however, the measurements are still corresponding to the qudit Pauli operators $T_{u}$. The expectation values have the same form as for the full Bell operator (16), but the summation and the maximum take only coefficients $x,z,t\in\mathbb{Z}_{d}$. The Bell operator is a measure of the singlet fraction and the volume of Wigner negativity. The Bell inequality is structurally the same and Bell state achieves the same Bell violation.
As a result, we construct a family of Bell inequalities based on the Wigner function of the Bell state rotated by cube unitary operators. In Appendix B, we outline how these can be regarded as a natural generalization of the Clauser-Horne-Shimony-Holt (CHSH) inequality.
@@
where $X_{(1/3)}=UXU^{\dagger}$, and $U=\mathrm{diag}(1,\omega^{2/3},\omega^{1/3})$. Note that these operators $O=X,X_{(1/3)}$ are not Hermitian but unitary yet represent valid measurements in terms of their Fourier transform with projectors $\Pi_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}O^{k}$ of outcome $m\in\mathbb{Z}_{d}$. The same holds for a deterministic value assignment. As a result, any lhv model attains $B_{3,\mathrm{lhv}}^{\max}=\max(\omega^{a_{0}+b_{0}}+\omega^{2a_{0}+2b_{1}}+\omega^{2a_{1}+2b_{0}}+\omega^{a_{1}+b_{1}+1}+\,h.c.)=6+\omega+\omega^{2}=5$, where the maximum is taken over $a_{0},b_{0},a_{1},b_{1}\in\mathbb{Z}_{3}$, and is attained for $b_{1}=2$ and $a_{0},b_{0},a_{1}=0$. On the other hand, the quantum expectation value, $\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)=1+3(2\omega^{1/3}+\omega^{-2/3})+\,h.c.\,\approx 5.412\,,$ exceeds that of the classical model, $B_{3,\mathrm{lhv}}^{\max}<\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)$.
@@
$\mathcal{B}_{(1/2)}$ $=\frac{1}{d}\sum_{k\in\mathbb{Z}_{d}}X^{k}_{(1/2)}\otimes\left(X^{k}_{(-1/2)}+\omega^{-k}X^{k}_{(1/2)}\right)$
$+X^{k}_{(-1/2)}\otimes\left(X^{k}_{(1/2)}+\omega^{k}X^{k}_{(-1/2)}\right)\,.$ (23)
@@
The operator $\mathcal{B}_{S}$ is a measure for the overlap with the given stabilizer state under the condition that $((0,\mathbb{Z}_{d})_{1}\otimes(0,0)^{\otimes n-1})\subset\Sigma$, which is exactly the case if $|S\rangle$ is entangled over the cut of the first qudit. Then, $\langle S|\mathcal{B}_{S}|S\rangle=1$, while $\langle\mathcal{B}_{S}\rangle_{\mathrm{nc}}\leq d^{n}\max_{{\bm{u}}\in\Sigma}W_{{\bm{u}}}^{\nu}<1$, with the same arguments as for the bipartite case. Note that a single Bell operator of the form (24) cannot detect genuine multipartite entanglement GUHNE20091 beyond three parties.
@@
We have constructed several families of qudit Bell inequalities for any odd prime dimension. First, we propose a novel form of general Bell inequalities where a bipartite entangled state exhibiting Wigner negativity acts as a witness for nonlocality. Then, we construct a family of Bell operators aiming at a nonlocal violation with the Bell state. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $T$-gate characterized in Howard2012 . In Eq. (29), we illustrate how the Bell inequalities can be regarded as a natural extension of the CHSH inequality. The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. The maximally entangled Bell state violates the corresponding Bell inequality and attains the quantum bound due to the separation between the 1-norm and the maximum norm of its rotated Wigner function. Furthermore, the Bell operator measures the singlet fraction and bounds the extent of Wigner negativity. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. The measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models. Moreover, nonlocal correlations can emerge with qudit Pauli operators when rotated by unitary operators with a non-character spectrum in bipartite systems. These operators have an obvious advantage, which is resorting to phases $\omega^{q}$ that are beyond the description of any local value assignment that can only resort to characters $\omega^{k}$. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving GKP (Gottesman–Kitaev–Preskill) states PhysRevA.64.012310 .
@@ Appendix A Wigner Negativity
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}j,k\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\omega^{[u_{1},v_{1}]+[u_{2},v_{2}]+k((v_{1})_{z}+(v_{2})_{z})+2^{-1}((v_{1})_{x}(v_{1})_{z}+(v_{2})_{x}(v_{2})_{z})}\langle jj|U^{\phantom{\dagger}}_{{f}}\,X^{(v_{1})_{x}}U^{\dagger}_{{f}}\otimes X^{(v_{2})_{x}}|kk\rangle$
@@ Appendix B Relation to the CHSH inequality
with the Pauli operators $\sigma_{(0,0)}=\mathds{1}$, $\sigma_{(1,0)}=X$, $\sigma_{(1,1)}=Y$, $\sigma_{(0,1)}=Z$. The evaluations of the expectation values and nonlocal violations for the qudit Bell operators in Eq. (16) and Eq. (20) do not straightforwardly generalize to qubits using the above Wigner function. The reason is that the assignments of an lhv model are different from qudit Pauli operators to qubit Pauli operators if one measures a complete set of operators generating mutually unbiased bases. However, in case where $U_{\nu}=T$, the equivalent Bell operator for qubits in Eq. (16) is
$\mathcal{B}^{\phantom{\prime}}_{2}=\mathds{1}\otimes\mathds{1}+\sigma^{z}\otimes\sigma^{z}-\frac{1}{\sqrt{2}}\Big{(}T\sigma^{y}T^{\dagger}\otimes\sigma^{y}+T\sigma^{x}T^{\dagger}\otimes\sigma^{y}+T\sigma^{y}T^{\dagger}\otimes\sigma^{x}-T\sigma^{x}T^{\dagger}\otimes\sigma^{x}\Big{)}\,,$ (29)
leading to an inequality $1+A_{2}B_{2}-\left(A_{0}B_{0}+A_{1}B_{0}+A_{0}B_{1}-A_{1}B_{1}\right)/\sqrt{2}\leq 2+\sqrt{2}$ that is analogous to the Clauser-Horne-Shimony-Holt (CHSH) $A_{0}B_{0}+A_{1}B_{0}+A_{0}B_{1}-A_{1}B_{1}\leq 2$.
Appendix C Bounds on Gross’ Wigner function
We consider a general pure bipartite qudit state $|\Psi\rangle$. The sum over absolute values can be expressed by the negative sum of all negative values, the negativity volume $N[\Psi]$, with $\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}|W_{v_{1},v_{2}}[\Psi]|=1+2N[\Psi]$. Hölder’s inequality bounds the negativity volume through
$\Bigg{(}\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}\left|W_{v_{1},v_{2}}[\Psi]\right|\,\Bigg{)}^{2}\leq d^{4}\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W_{v_{1},v_{2}}[\Psi]^{2}=d^{2}\,.$ (30)
Altogether, it is
$\frac{1}{d^{3}}\leq\frac{1}{d^{2}\left(1+2N[\Psi]\right)}\leq\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}|W_{v_{1},v_{2}}[\Psi]|\leq\frac{1}{d}\,,$ (31)
for any pure bipartite state $|\Psi\rangle$.
@@ Appendix C Bounds on Gross’ Wigner function
$\mathrm{Tr}\left(\sigma\mathcal{B}\right)$ $=d^{2}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{u_{1},u_{2}}W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\sigma\,(U_{\nu}\otimes\mathds{1})\right)$
$\leq d^{2}\max_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{u_{1},u_{2}}|\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\sigma\,(U_{\nu}\otimes\mathds{1})\right)|\,.$ (32)
Appendix D Bell Inequality for Stabilizer States